What is the nth term rule of the linear sequence below?


27

,

25

,

23

,

21

,

19

,



1
Expert's answer
2022-01-31T16:13:30-0500

The linear sequence 27,25,23,21,19,... is an arithmetic progression with a common difference of -2. For an arithmetic progression, the "n^{th}" term is given as, "a_n=a_1+(n-1)d"  where "a_n" is the "n^{th}" term, "d=-2" is the common difference and "a_1=27" is the first term.

So, the "n^{th}" term is given as, "a_n=27-2(n-1)=27-2n+2=29-2n"

Therefore the n"^{th}" term rule is "a_n=29-2n".